OpenStudy (anonymous):

The Johnson twins were born 11 years after their older sister. This year, the product of the three siblings ages is exactly 583 more than the sum of their ages. How old are the twins?

OpenStudy (anonymous):

Let x = the age of their older sister and y = age of each twin. 1) y = x+11 2) xy^2 = (x+2y) + 583 You know what y equals (x+11) so plug into the second equation to get an equation with only x's. Once you solve for x, put the value back into equation 1 to get the age of the twins.